# Amnesiac Flooding: Easy to break, hard to escape

### v1

@@ Abstract.
In this paper we make progress on both of these fronts. Under a reasonable restriction (obliviousness to message content) additional to the fault-free synchronous model, we prove that Amnesiac Flooding is the only strictly stateless deterministic protocol that can achieve terminating broadcast. We identify four natural properties of a terminating broadcast protocol that Amnesiac Flooding uniquely satisfies. In contrast, we prove that even minor relaxations of any of these four criteria allow the construction of other terminating broadcast protocols.
@@
ccs: Mathematics of computing Graph algorithmsccs: Mathematics of computing Discrete mathematicsccs: Theory of computation Graph algorithms analysisccs: Theory of computation Distributed algorithms
@@ 1. Introduction
The dissemination of information to disparate participants is a fundamental problem in both the construction and theory of distributed systems. A common strategy for solving this problem is to “broadcast”, i.e. to transmit a piece of information initially held by one agent to all other agents in the system. In fact, broadcast is not merely a fundamental communication primitive in many models, but also underlies solutions to other fundamental problems such as leader election and wake-up. Given this essential role in the operation of distributed computer systems and the potential volume of broadcasts, an important consideration is simplifying the algorithms and minimizing the overhead required for each broadcast.
@@
Amnesiac flooding algorithm. (adapted from (Hussak and Trehan, 2023)) Let $G=(V,E)$ be an undirected graph, with vertices $V$ and edges $E$ (representing a network where the vertices represent the nodes of the network and edges represent the connections between the nodes). Computation proceeds in synchronous ‘rounds’ where each round consists of nodes receiving messages sent from their neighbours. A receiving node then sends messages to some neighbours in the next round. No messages are lost in transit. The algorithm is defined by the following rules:
@@
Outside of a partial robustness to crash failures, the fault sensitivity of Amnesiac Flooding under synchrony has not been explored in the literature. This omission is further compounded by the use of Amnesiac Flooding as an underlying subroutine for the construction of other broadcast protocols. In particular, multiple attempts have been made to extend Amnesiac Flooding to new settings (for example routing multiple concurrent broadcasts (Bayramzadeh et al., 2021) or flooding networks without guaranteed edge availability (Turau, 2021b)), while maintaining its desirable properties. However, none have been entirely successful, typically requiring some state-fulness. It has not in fact been established that any other protocol can retain all of Amnesiac Flooding’s remarkable properties even in its original setting. These gaps stem fundamentally from the currently limited knowledge of the dynamics of Amnesiac Flooding beyond the fact of its termination and its speed to do so. In particular, both of the existing techniques (parity arguments such as in (Hussak and Trehan, 2023) or auxiliary graph constructions such as in (Turau, 2021a)) used to obtain termination results for Amnesiac Flooding are unable to consider faulty executions of the protocol and fail to capture the underlying structures driving terminating behaviour.
We address these gaps through the application of novel analysis and by considering the structural properties of Amnesiac Flooding directly. By considering the sequence of message configurations, we are able to identify the structures underlying Amnesiac Flooding’s termination and use these to reason about the algorithm in arbitrary configurations. The resulting dichotomy gives a comprehensive and structured understanding of termination in Amnesiac Flooding. For example, we apply this to investigate the sensitivity of Amnesiac Flooding with respect to several forms of fault and find it to be quite fragile. Furthermore, we show that under reasonable assumptions on the properties of a synchronous network, any strictly stateless deterministic terminating broadcast algorithm oblivious to the content of messages, must produce the exact same sequence of message configurations as Amnesiac Flooding on any network from any initiator. We therefore argue that Amnesiac Flooding is unique. However, we show that if any of these restrictions are relaxed, even slightly, distinct terminating broadcast algorithms can be obtained. As a result of this uniqueness and simplicity, we argue that Amnesiac Flooding represents a prototypical broadcast algorithm. This leaves open the natural question: do there exist fundamental stateless algorithms underlying solutions to other canonical distributed network problems?
@@ 1.1. Our Contributions
Our main technical results regarding the existence of alternative protocols to Amnesiac Flooding are given in the next two theorems (reworded in Sections 4 and 5).
Theorem 0 (Uniqueness of Amnesiac Flooding).
@@
Note that the last theorem allows, but does not require, that nodes have access to unique identifiers labelling themselves and their ports. However, we enforce the condition that these identifiers, should they exist, may be drawn adversarially from some super set of $[n+\kappa]$ where $n$ is the number of nodes on the networks and $\kappa=R(9,8)$ where $R(9,8)$ is a Ramsey Number. It is important to stress here that this result holds even if the space of unique identifiers is only greater than $n$ by an additive constant. In contrast to the last theorem, the next one does not assume that agents have access to unique identifiers.
Theorem 0 (Existence of relaxed Algorithms).
@@
    Neighbourhood-2 Flooding: Strict Statelessness  is relaxed.
@@
    Random Flooding: Determinism  is relaxed.
@@
    1-Bit Flooding (Message Dependent): Obliviousness  is relaxed.
@@
    1-Bit Flooding (High Bandwidth): Unit Bandwidth is relaxed.
@@
The main technical result concerning fault sensitivity is a dichotomy characterizing the configurations, starting from which, Amnesiac Flooding terminates. As the rigorous statement of the result requires some additional notation and terminology, we will state it only informally here. We show in Theorem 6.1 that, whether a configuration terminates or not, is determined exclusively by the parity of messages distributed around cycles and the so-called faux-even cycles (FEC). This result implies the following three theorems which demonstrate the fragility of Amnesiac Flooding under three increasingly strong forms of fault. We give our fault model explicitly in Section 6.1.
Theorem 0 (Single Message Failure).
If single-source Amnesiac Flooding experiences a single message drop failure for the message $(u,v)$ then it fails to terminate. if either:
  • $uv$ is not a bridge
  • $uv$ lies on a FEC subgraph
Moreover, it fails to broadcast if and only if this is the first message sent along $uv$, $uv$ is a bridge, and the side of the cut containing $u$ does not contain an odd cycle.
We, also, consider the possibility of a link/edge failing in one direction (or if we consider an undirected edge as two directed links in opposite directions, only one of the directed links fails).
Theorem 0 (Uni-directional link failure).
@@
Theorem 0 (Byzantine Failure).
@@
  • Non-termination if and only if at least one member of $J$ lies on either a cycle or FEC subgraph.
@@
Organisation of the paper: The initial part of the paper presents the required technical and motivational background, statements of our results and a technical outline of the more interesting proofs. The detailed proofs are presented in the appendix. Related work is presented in Section 2. Section 3 presents the model and notation required for the following technical sections. Uniqueness of the algorithm is discussed in Section 4 with full proofs given in appendix B. We relax the conditions individually to derive some more stateless protocols in Section 5 with the detailed proofs in appendix C. We present our work on the fault sensitivity of amnesiac flooding in Section 6 with the detailed proofs in appendix D. However, for technical reasons, we give the full proof of  Theorem 6.1 earlier in appendix A. We end with our conclusions and pointers to future work in Section 7.
@@ 2. Related Work
The termination of Amnesiac Flooding and its derivatives has been the focus of several works. Two independent proofs of the algorithm’s termination have been presented, using either parity arguments over message return times (Hussak and Trehan, 2023) or auxillary graph constructions (Turau, 2021a). The latter technique has further been used to establish tight diameter independent bounds on the termination of multicast using Amnesiac Flooding, complementing the eccentricity based bounds of (Hussak and Trehan, 2023). The techniques we develop in this work combine notions of both and allows us to consider arbitrary configurations of messages, rather than those resulting purely from correct broadcast.
@@
There have, further, been multiple variants of the algorithm introduced. It was observed by (Turau, 2020) in a result reminiscent of the BASIC protocol proposed by (Gopal et al., 1999), that sending a second wave of messages from a subset of the initial nodes could reduce the worst case $12D+$ 1 termination time to the optimal $D$ in all but a specific subset of bipartite graphs. We note that our fault sensitivity results extend naturally to this algorithm as well, as the same invariants apply to this setting. Beyond this, there have been several approaches to deal with the flooding of multiple messages simultaneously. In (Hussak and Trehan, 2020b), the original authors of (Hussak and Trehan, 2023) show that under certain conditions termination can be retained, even when conflicting floods occur. Since then two partially stateless algorithms have been proposed, both making use of message buffers and a small amount of local memory (Turau, 2021b; Bayramzadeh et al., 2021). We will not be directly concerned with these approaches, however, as we assume a single concurrent broadcast throughout. However, the mechanism employed in (Turau, 2021b) should be highlighted as it rather cleverly exploits the underlying parity properties we identify as driving termination. Furthermore, as reduction to Amnesiac Flooding is used as a technique for proofs in many of these works, the comprehensive understanding of its termination we present here could prove a powerful tool for future work in these areas.
@@
While the robustness of Amnesiac flooding and its variants have been previously studied, this has been focused on two forms of fault. The first is the disappearance and reappearance of nodes and links. The termination of Amnesiac Flooding is robust to the former and vulnerable to the latter (Hussak and Trehan, 2023). We will observe that this is a necessary consequence of the invariants driving termination and their relation to cyclicity. A pseudo-stateless extension to Amnesiac Flooding has been proposed to circumvent this (Turau, 2021b), implicitly exploiting the parity conditions of (Hussak and Trehan, 2023). The second are faults that violate synchrony. Under a strong form of asynchrony, truly stateless and terminating broadcast is impossible (Turau, 2020). However, the landscape under a weaker form of asynchrony (namely, the case of fixed delays on communication links) is more fine-grained. Although termination results have been obtained for cycles, as well as the case of single delayed edges in bipartite graphs (Hussak and Trehan, 2023), there is no clear understanding of the impact of fixed channel delays. While we do not directly address this, we believe that techniques mirroring our invariant characterizations may prove fruitful in this area. To our knowledge, this work is the first to consider both the uniqueness of Amnesiac Flooding, as well as its fault sensitivity beyond node/link unavailability in a synchronous setting.
@@
Beyond Amnesiac Flooding and its extensions, the role of memory in information dissemination is well studied in a variety of contexts. Frequently, stateful methods obtain faster termination time, such as in the phone-call model where the ability to remember one’s communication partners and prevent re-communication dramatically improves termination time and message efficiency (Elsässer and Sauerwald, 2008; Berenbrink et al., 2014; Doerr et al., 2011). Similarly, for bit dissemination in the passive communication model the addition of only $\log{\log{n}}$ bits of memory is sufficient to break the near linear time convergence lower bound of (D’Archivio and Vacus, 2024) and achieve polylogarithmic time (Korman and Vacus, 2022). Even more strongly, a recent work (Parzych and Daymude, 2024) has shown that in the context of synchronous anonymous dynamic networks, stabilizing broadcast from an idle start is impossible with $O(1)$ memory and even with $o(\log{n})$ memory if termination detection is required. Despite this low memory and even stateless broadcasts remain desirable (Gopal et al., 1999). The possibility of solving other canonical distributed computing problems beyond broadcast, in a stateless manner, remains intriguing. In this direction, various low memory stateful models have been proposed to handle more complex distributed problems e.g. the compact local streaming (CLS) model in (Castañeda et al., 2020) with deterministic solutions (routing, self-healing fault-tolerance etc) and randomised solutions for distributed coloring in a similar model (Flin et al., 2023).
@@
Several stateless broadcast schemes exist, to give an example, for mobile ad hoc networks which due to the lower power and rapid movement of devices see diminishing returns from maintaining information about the network (Manfredi et al., 2011). However, given a lack of synchronisation as well as the wish to avoid so-called broadcast storms (Ni et al., 1999), these techniques typically rely on either some form of global knowledge (such as the direction or distance to the initiator) or the ability to sample network properties by eavesdropping on communications over time (Banerjee et al., 2012; Ruiz and Bouvry, 2015). It should be noted however, that in contrast to many models, such as anonymous dynamic networks, radio networks and many manets, the typical framework for studying stateless flooding (”true statelessness” as defined by (Turau, 2020) restricting the model of (Dolev et al., 2017)) permits the knowledge and distinguishing of neighbours in both broadcasting and receiving.

### v2

@@ Abstract.
In this paper we make progress on both of these fronts. Under a reasonable restriction (obliviousness to message content) additional to the fault-free synchronous model, we prove that Amnesiac Flooding is the only strictly stateless deterministic protocol that can achieve terminating broadcast. We achieve this by identifying four natural properties of a terminating broadcast protocol that Amnesiac Flooding uniquely satisfies. In contrast, we prove that even minor relaxations of any of these four criteria allow the construction of other terminating broadcast protocols.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Mathematics of computing Graph algorithmsccs: Mathematics of computing Discrete mathematicsccs: Theory of computation Graph algorithms analysisccs: Theory of computation Distributed algorithms
@@ 1. Introduction
The dissemination of information to disparate participants is a fundamental problem in both the construction and theory of distributed systems. A common strategy for solving this problem is to “broadcast”, i.e. to transmit a piece of information initially held by one agent to all other agents in the system (Attiya and Welch, 2004; Lynch, 1996; Peleg, 2000; Tanenbaum, 2011; Tel, 1994). In fact, broadcast is not merely a fundamental communication primitive in many models, but also underlies solutions to other fundamental problems such as leader election and wake-up. Given this essential role in the operation of distributed computer systems and the potential volume of broadcasts, an important consideration is simplifying the algorithms and minimizing the overhead required for each broadcast.
@@
Amnesiac flooding algorithm. (adapted from (Hussak and Trehan, 2023)) Let $G=(V,E)$ be an undirected graph, with vertices $V$ and edges $E$ (representing a network where the vertices represent the nodes of the network and edges represent the connections between the nodes). Computation proceeds in synchronous ‘rounds’ where each round consists of nodes receiving messages sent from their neighbours. A receiving node then sends messages to some neighbours which arrive in the next round. No messages are lost in transit. The algorithm is defined by the following rules:
@@
Outside of a partial robustness to crash failures, the fault sensitivity of Amnesiac Flooding under synchrony has not been explored in the literature. This omission is further compounded by the use of Amnesiac Flooding as an underlying subroutine for the construction of other broadcast protocols. In particular, multiple attempts have been made to extend Amnesiac Flooding to new settings (for example routing multiple concurrent broadcasts (Bayramzadeh et al., 2021) or flooding networks without guaranteed edge availability (Turau, 2021b)), while maintaining its desirable properties. However, none have been entirely successful, typically requiring some state-fulness. It has not in fact been established that any other protocol can retain all of Amnesiac Flooding’s remarkable properties even in its original setting. These gaps stem fundamentally from the currently limited knowledge of the dynamics of Amnesiac Flooding beyond the fact of its termination and its speed to do so. In particular, all of the existing techniques (e.g. parity arguments such as in (Hussak and Trehan, 2023) and auxiliary graph constructions such as in (Turau, 2021a)) used to obtain termination results for Amnesiac Flooding are unable to consider faulty executions of the protocol and fail to capture the underlying structures driving terminating behaviour.
We address these gaps through the application of novel analysis and by considering the structural properties of Amnesiac Flooding directly. By considering the sequence of message configurations, we are able to identify the structures underlying Amnesiac Flooding’s termination and use these to reason about the algorithm in arbitrary configurations. The resulting dichotomy gives a comprehensive and structured understanding of termination in Amnesiac Flooding. For example, we apply this to investigate the sensitivity of Amnesiac Flooding with respect to several forms of fault and find it to be quite fragile. Furthermore, we show that under reasonable assumptions on the properties of a synchronous network, any strictly stateless deterministic terminating broadcast algorithm oblivious to the content of messages, must produce the exact same sequence of message configurations as Amnesiac Flooding on any network from any initiator. We therefore argue that Amnesiac Flooding is unique. However, we show that if any of these restrictions are relaxed, even slightly, distinct terminating broadcast algorithms can be obtained. As a result of this uniqueness and simplicity, we argue that Amnesiac Flooding represents a prototypical broadcast algorithm. This leaves open the natural question: do there exist fundamental stateless algorithms underlying solutions to other canonical distributed network problems? Though memory can be essential or naturally useful in certain scenarios (Parzych and Daymude, 2024; D’Archivio and Vacus, 2024; Korman and Vacus, 2022; Elsässer and Sauerwald, 2008; Berenbrink et al., 2014; Doerr et al., 2011), understanding what we can do with statelessness can help us push fundamental boundaries.
@@ 1.1. Our Contributions
Our main technical results regarding the existence of alternative protocols to Amnesiac Flooding are given in the next two theorems (reworded in Sections 4 and 4.2).
Theorem 1.2 (Uniqueness of Amnesiac Flooding).
@@
Note that the last theorem allows, but does not require, that nodes have access to unique identifiers labelling themselves and their ports. However, we enforce the condition that these identifiers, should they exist, may be drawn adversarially from some super set of $n+\kappa]$ where $n$ is the number of nodes on the networks and $\kappa=R(9,8)$ where $R(9,8)$ is a Ramsey Number. It is important to stress here that this result holds even if the space of unique identifiers is only greater than $n$ by an additive constant. In contrast to the last theorem, the next one does not assume that agents have access to unique identifiers.
Theorem 1.3 (Existence of relaxed Algorithms).
@@
    Neighbourhood-2 Flooding: Strict Statelessness is relaxed and agents are given knowledge of their neighbours’ neighbours. This allows for distinct behaviour on networks of radius one as some agents are aware of the entire network topology.
@@
    Random Flooding: Determinism is relaxed and agents are given access to a random coin. Agents randomly choose in each round whether to use Amnesiac Flooding or to forward messages to all of their neighbours.
@@
    1-Bit Flooding (Message Dependent): Obliviousness is relaxed and the source is allowed to include one bit of read-only control information in the messages. The source records in the message whether they are a leaf vertex, and if so agents perform Amnesiac Flooding upon receiving the message. Otherwise, agents perform a modified version where leaf vertices return messages.
@@
    1-Bit Flooding (High Bandwidth): Unit Bandwidth is relaxed and agents are allowed to send either one or two messages over an edge. This permits the same algorithm as the previous case by encoding the control information in the number of messages sent.
@@
The main technical result concerning fault sensitivity is a dichotomy characterizing the configurations, starting from which, Amnesiac Flooding terminates. As the rigorous statement of the result requires some additional notation and terminology, we will state it only informally here. We show in Theorem 5.1 that, whether Amnesiac Flooding terminates when begun from a configuration or not, is determined exclusively by the parity of messages distributed around cycles and so-called faux-even cycles (FEC) (essentially pairs of disjoint odd cycles connected by a path, see section 5 for a full definition) within the graph. This result implies the following three theorems which demonstrate the fragility of Amnesiac Flooding under three increasingly strong forms of fault. We give our fault model explicitly in Section 5.2.
Theorem 1.4 (Single Message Failure).
If single-source Amnesiac Flooding experiences a single message drop failure for the message $(u,v)$ then it fails to terminate if either:
  • uvuv is not a bridge
  • uvuv lies on a path between odd cycles
Moreover, it fails to broadcast if and only if this is the first message sent along $uv$, uvuv is a bridge, and the side of the cut containing $u$ does not contain an odd cycle.
We also consider the possibility of a link/edge failing in one direction (or if we consider an undirected edge as two directed links in opposite directions, only one of the directed links fails).
Theorem 1.5 (Uni-directional link failure).
@@
Theorem 1.6 (Byzantine Failure).
@@
  • Non-termination if and only if at least one member of $J$ lies on either a cycle or a path between odd-cycles.
@@
Organisation of the paper: The initial part of the paper presents the required technical and motivational background, statements of our results and a technical outline of the more interesting proofs. The detailed proofs are presented in the appendix. Related work is presented in Section 2. Section 3 presents the model and notation required for the following technical sections. Uniqueness of the algorithm is discussed in Section 4 with full proofs given in appendix B. We relax the conditions individually to derive some more stateless protocols in Section 4.2 with the detailed proofs in appendix C. We present our work on the fault sensitivity of amnesiac flooding in Section 5 with the detailed proofs in appendix D. However, for technical reasons, we give the full proof of  Theorem 5.1 earlier in appendix A. We end with our conclusions and pointers to future work in Section 6.
@@ 2. Related Work
The termination of Amnesiac Flooding and its derivatives has been the focus of several works. In combination they provide the following result that Amnesiac Flooding terminates under any sequence of multi-casts, i.e.
Theorem 2.1 (Termination of Amnesiac Flooding (adapted from (Hussak and Trehan, 2020b, 2023; Turau, 2021a))).
For $G=(V,E)$ a graph and $I_{1},...,I_{k}\subseteq V$ a sequence of sets initiator nodes, Amnesiac Flooding on $G$ terminates when initiated from $I_{1}$ in round $1$, then $I_{2}$ in round $2$ and so forth.
@@
Two independent proofs of the algorithm’s termination have been presented, using either parity arguments over message return times (Hussak and Trehan, 2023) or axillary graph constructions (Turau, 2021a). The latter technique has further been used to establish tight diameter independent bounds on the termination of multicast using Amnesiac Flooding, complementing the eccentricity based bounds of (Hussak and Trehan, 2023). The techniques we develop in this work, however, are more closely aligned with those of (Hussak and Trehan, 2020b), as we exploit a similar notion to their ”even flooding cycles” in our message path argument. In contrast to that work we focus on arbitrary configurations of messages, rather than just those resulting purely from a correct broadcast which allows us to obtain a much stronger characterisation. Combining our techniques with the dual ”reverse” flood introduced by (Hussak and Trehan, 2020b), we are able to show the complement of Theorem 2.1, that only those configurations reachable from a sequence of multi-casts lead to termination.
@@
There have, further, been multiple variants of Amnesiac Flooding introduced. It was observed by (Turau, 2020) in a result reminiscent of the BASIC protocol proposed by (Gopal et al., 1999), that sending a second wave of messages from a subset of the initial nodes could reduce the worst case $2D+1$ termination time to the optimal $D+1$ in all but a specific subset of bipartite graphs. We note that our fault sensitivity results extend naturally to this algorithm as well, as the same invariants apply to this setting. Beyond this, there have been several approaches to deal with the flooding of multiple messages simultaneously. In (Hussak and Trehan, 2020b), the original authors of (Hussak and Trehan, 2019) show that under certain conditions termination can be retained, even when conflicting floods occur. Since then, two partially stateless algorithms have been proposed, both making use of message buffers and a small amount of local memory (Turau, 2021b; Bayramzadeh et al., 2021). We will not be directly concerned with these approaches, however, as we assume a single concurrent broadcast throughout. However, the mechanism employed in (Turau, 2021b) should be highlighted as it rather cleverly exploits the underlying parity properties we identify as driving termination. Furthermore, as reduction to Amnesiac Flooding is used as a technique for proofs in many of these works, the comprehensive understanding of its termination we present here could prove a powerful tool for future work in these areas.
@@
While the robustness of Amnesiac flooding and its variants have been previously studied, this has been focused on two forms of fault. The first is the disappearance and reappearance of nodes and links. The termination of Amnesiac Flooding is robust to disappearance and vulnerable to reappearance (Hussak and Trehan, 2023). We will observe that this is a necessary consequence of the invariants driving termination and their relation to cyclicity. In particular, the disappearance of nodes and links cannot form new cycles violating the invariant, whereas their reappearance can. A pseudo-stateless extension to Amnesiac Flooding has been proposed to circumvent this (Turau, 2021b), implicitly exploiting the parity conditions of (Hussak and Trehan, 2023). The second are faults that violate synchrony. Under a strong form of asynchrony, truly stateless and terminating broadcast is impossible (Turau, 2020). However, the landscape under a weaker form of asynchrony (namely, the case of fixed delays on communication links) is more fine-grained. Although termination results have been obtained for cycles, as well as the case of single delayed edges in bipartite graphs (Hussak and Trehan, 2023), there is no clear understanding of the impact of fixed channel delays. While we do not directly address this, we believe that techniques mirroring our invariant characterizations may prove fruitful in this area. To our knowledge, this work is the first to consider both the uniqueness of Amnesiac Flooding, as well as its fault sensitivity beyond node/link unavailability in a synchronous setting.
@@
Beyond Amnesiac Flooding and its extensions, the role of memory in information dissemination is well studied in a variety of contexts. Frequently, stateful methods obtain faster termination time, such as in the phone-call model where the ability to remember one’s communication partners and prevent re-communication dramatically improves termination time and message efficiency (Elsässer and Sauerwald, 2008; Berenbrink et al., 2014; Doerr et al., 2011). Similarly, for bit dissemination in the passive communication model the addition of only $\log{\log{n}}$ bits of memory is sufficient to break the near linear time convergence lower bound of (D’Archivio and Vacus, 2024) and achieve polylogarithmic time (Korman and Vacus, 2022). Even more strongly, a recent work (Parzych and Daymude, 2024) has shown that in the context of synchronous anonymous dynamic networks, stabilizing broadcast from an idle start is impossible with $O(1)$ memory and even with $o(\log{n})$ memory if termination detection is required. Despite this, low memory and even stateless broadcasts remain desirable (Gopal et al., 1999). The possibility of solving other canonical distributed computing problems beyond broadcast, in a stateless manner, remains intriguing. In this direction, various low memory stateful models have been proposed to handle more complex distributed problems e.g. the compact local streaming (CLS) model in (Castañeda et al., 2020) with deterministic solutions (routing, self-healing fault-tolerance etc.) and randomised solutions for distributed colouring in a similar model (Flin et al., 2023).
@@
Several stateless broadcast schemes exist. To give an example, they are used in mobile ad hoc networks, which, due to the lower power and rapid movement of devices, see diminishing returns from maintaining information about the network (Manfredi et al., 2011). However, given a lack of synchronisation as well as the wish to avoid so-called broadcast storms (Ni et al., 1999), these techniques typically rely on either some form of global knowledge (such as the direction or distance to the initiator) or the ability to sample network properties by eavesdropping on communications over time (Banerjee et al., 2012; Ruiz and Bouvry, 2015). It should be noted however, that in contrast to many models, such as anonymous dynamic networks, radio networks and many manets, the typical framework for studying stateless flooding (”true statelessness” as defined by (Turau, 2020) restricting the model of (Dolev et al., 2017)) permits the knowledge and distinguishing of neighbours in both broadcasting and receiving.